# Synthetic Prior for Few-Shot Drivable Head Avatar Inversion

### v1

@@ Acknowledgement
The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting WZ. JT is supported by the Google Gift Award.
@@ Appendix A Appendix
All our inversion results used only three input images unless stated otherwise. Figure 19 compares monocular baseline methods trained on the entire dataset with our inversion approach. Furthermore, we provide additional examples of cross-reenactment comparisons, demonstrating the advantages of our method compared to baselines trained on only 13 frames. Next, we present results with progressively varying numbers of training frames, illustrating how this influences the quality of reconstruction. Figures 20, 21, and 22 highlight the importance of our synthetic prior.

### v2

@@ Acknowledgement
The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting WZ. JT was supported by the ERC Starting Grant LeMo (101162081). We also would like to thank Daoye Wang for assisting with synthetic asset generation, Mark Murphy for his help with using Google infrastructure, and Menglei Chai for providing the hair proxy.
@@ Appendix A Appendix
All our inversion results used only three input images (Figure 22) unless stated otherwise. Figure 17 compares monocular baseline methods trained on the entire dataset with our inversion approach. Furthermore, we provide additional examples of cross-reenactment comparisons, demonstrating the advantages of our method compared to baselines trained on only 13 frames. Next, we present results with progressively varying numbers of training frames, illustrating how this influences the quality of reconstruction. Figures 18 and 19 highlight the importance of our synthetic prior.
@@ Appendix C Broader Impact
Figure 22: Unless otherwise stated, all experiments in this paper used three input images. Here, we present these images for each actor.